What is 3(x-2)-2(x+7)=2x+17 step by step?
Found 2 solutions by tracybaker_1, mananth:
Answer by tracybaker_1(1)

What is 3(x-2)-2(x+7)=2x+17 step by step?
3x-6-2x+14=2x+17
x+8=2x+17
x+17=8
x+17-17=8-17
x=-9

Answer by mananth(12269)

3(x-2)-2(x+7)=2x+17 step by step?
3x-6-2x-14=2x+17
regroup the terms
3x-2x-6-14=2x+17
x-20=2x+17
add -x to both sides
x-x-20=2x-x+17
-20=x+17
add -17 to both sides
-20-17 =x+17-17
-37=x